OpenStudy (anonymous):

do any one know how to identify the solutions of equation 2x+y=1 and x-y=3 if any

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

2x+y=1 x-y=4 + --------------- 3x+0y=5 3x=5 x=5/3 you can find y by substitution.

OpenStudy (anonymous):

Hi, can you give in more details how you receive 5/3 as your answer

OpenStudy (anonymous):

apparently i am wrong

OpenStudy (anonymous):

oh i made a typo let me do that again

OpenStudy (anonymous):

2x+y=1 x-y=3 + ------------- 3x+0y=4 3x=4 x= 4/3

OpenStudy (anonymous):

now it's correct. I'll explain the steps

OpenStudy (anonymous):

To find the solution, you want to eliminate either x or y, so that you only have 1 unknown instead of 2. Not all the equations are the same, but in this one because you have a y in one equation and a -y in the other it's the easiest to add them together so that you will end up with 0y (so only an x)

OpenStudy (anonymous):

in other cases it's better to subtract, and sometimes you have to multiply one of the equations before you proceed. If you want i can give some examples on that later.

OpenStudy (anonymous):

by the way two lines ALWAYS cross once, EXCEPT if the lines have the same slope (then they will either NEVER cross or ALWAYS cross, if they are on top of each other)

OpenStudy (anonymous):

hey you don't think the possible solution could be both 4/3,-5/3

OpenStudy (anonymous):

that's the x and y solution, you find the y after you have found the x

OpenStudy (anonymous):

if you have found the value for x, you plug the x value in one of the two original equations and you'll automatically find y

OpenStudy (anonymous):

how is that done and also try identifying the solutions of this equation if there's any 3x-2y=6 and 5x=5-25y

OpenStudy (anonymous):

well, since you know x=4/3 then you take 2x+y=1 and fill in x so that 2*(4/3)+y=1 and solve for y

OpenStudy (anonymous):

3x-2y=6 and 5x=5-25y rewrite to 3x-2y=6 and 5x+25y=5 multiply so both equations have the same amount of x in it. 3x-2y=6 multiply the whole thing by 5 (to get 15x) 5x+25y=5 multiply the whole thing by 3 (to get 15x) 15x-10y=30 15x+75y=15 subtract them from each other and proceed. I'll get something to eat now, i'll be back later and see if you were able to solve it.